# CVOCSemRPL: Class-Variance Optimized Clustering, Semantic Information Injection and Restricted Pseudo Labeling based Improved Semi-Supervised Few-Shot Learning

### v2

@@ 8.4 Effect of unlabeled samples
In this section, we analyze how our results vary with the availability of unlabeled samples. As we increase the number of unlabeled samples, we observe a corresponding improvement in performance, which is intuitive (see Tables 8 and 9). We also compare our method with Cluster-FSL [4], which represents the previous state-of-the-art architecture in semi-supervised few-shot learning prior to our work.
9 Ablation Experiments
9.1 Ablation on the contributions of the Cluster Separation Tuner and Semantic Correction
@@
In this section, in order to isolate the contributions of the Cluster Separation Tuner (CST) and semantic correction, we disable the reconstruction-distance term and plug CST + semantic anchor directly into standard clustering methods. Here, we compare against the latest methods under the 5-way 1-shot and 5-shot tasks. Although we had to retune a few hyper-parameters - chiefly $s$ in the semantic anchor module and the number of CST iterations to fit this plug-in setup, the adjusted settings still yield competitive performance (Figure 5).
9.2 Ablation on hyperparameter $s$ in Semantic anchor
@@
9.3 Ablation on Restricted Pseudo Labeling
At test time, we limit the augmentation of the support set by selecting only those unlabeled samples with the most confident pseudo-labels. Since pseudo-labels can be noisy, we sort unlabeled samples by their prediction entropy and pick the lowest-entropy examples, offering a simple, efficient strategy with minimal test-time overhead (see Figure 7). We observe that retaining 80% of the unlabeled samples yields the highest gains, highlighting the impact of confident pseudo-label selection.
@@ 9.3 Ablation on Restricted Pseudo Labeling
9.4 Choice of embedding model
We evaluate two text encoders for generating class-description embeddings: a vanilla BERT-base model [40] and the CLIP text encoder [27]. In both 5-way 1-shot and 5-way-5-shot settings, CLIP’s text features outperform BERT significantly (see Figure 8).
Potential reasons might be that CLIP’s text encoder is trained jointly with its image encoder on millions of image–caption pairs via a contrastive objective, grounding its representations in visual concepts. BERT, by contrast, learns only linguistic co-occurrence patterns and lacks alignment with vision. Consequently, CLIP’s text vectors naturally becomes far more effective option for bridging semantic gaps in few-shot vision tasks than a standalone language model.

### v3

@@ 8.4.5 Effect of unlabeled samples
In this section, we analyze how our results vary with the availability of unlabeled samples. As we increase the number of unlabeled samples, we observe a corresponding improvement in performance, which is intuitive (see Tables 12 and 13). We also compare our method with Cluster-FSL [ling2022semi], which represents the previous state-of-the-art architecture in semi-supervised few-shot learning prior to our work.
8.4.6 Quality of Representation Space and Prototype Refinement
Prior methods in SSFSL often commit to either global prototype-based structures (e.g., ProtoNet [snell2017prototypical], Soft K-Means [ren2018meta]) or reconstruction-based alignments (e.g., MFC [ling2022semi]). Prototype-based methods, while offering clear global decision boundaries, can falter in capturing fine-grained intra-class variations and local geometric nuances. Conversely, reconstruction-based clustering excels at modelling local manifolds but often mis-assigns samples when two class subspaces overlap, because they ignore where class centroids sit in the broader topology. We overcome this dichotomy by fusing the reconstruction based distance with explicit Euclidean regularizers: an intra-class term that contracts each cluster and an inter-class term that repels different centroids. A point is accepted by a class only if it is simultaneously well reconstructed by that class’s subspace and lies near its prototype while far from others. This fusion retains the fine-grained shape of each class manifold while simultaneously positioning the manifolds far enough apart to maintain clear global separation. The cluster separation tuner (CST) also adjusts each centroid to maximize intra-class compactness and inter-class separation. Applying this clustering method during the fine-tuning phase enhances the quality of the model’s learned representations.
In the testing phase, we nudge each visual prototype a small step toward the CLIP [radford2021learning] text embedding of its class description using our semantic injection network. This language-based anchor acts like a compass: it steers prototypes away from visually misleading regions (e.g., background clutter or occlusion) and pulls them toward positions that reflect the class’s true meaning. That is particularly valuable when we have only a handful of labeled examples. During inference, CVOC coupled with semantic anchoring enables the identification of superior class prototypes which is conceptually aligned with the class description. This enhanced discriminability yields substantially higher pseudo-label accuracy than prior approaches (see Table 5, where we compared our method against Kmeans and MFC [ling2022semi]), which in turn drives significant gains in FSL performance.
8.5 Ablation Experiments
8.5.1 Rationale for Pretraining and Finetuning
In our approach, pretraining learns a robust, general feature extractor from base classes, while finetuning adapts the model to few-shot episodes incorporating our novel components like CVOC. This clear separation ensures effective representation learning and task-specific adaptation. Empirically, skipping pretraining drops accuracy drastically (e.g., from 84.5% to 27% on MiniImageNet 5-way 1-shot with ResNet-12), and omitting finetuning also reduces performance significantly (to about 74.7%). These results highlight the essential roles of both stages in achieving strong few-shot performance.
8.5.2 Results on Inductive Setup
Although our primary experiments were conducted under the transductive setting, we also adapt the algorithm to evaluate it in the inductive setup during testing. The results presented in Table 14 demonstrate that our method maintains strong performance, highlighting its robustness in this scenario as well.
In the inductive setting, the model must classify each query sample independently using only the labeled support set, without accessing other unlabeled query samples during prediction. In contrast, the transductive setting allows the model to jointly leverage the entire unlabeled query set at test time, often enabling label propagation or graph-based refinement.
In short: inductive FSL predicts each query in isolation, while transductive FSL predicts them jointly using information from the whole query distribution.
Table 14: Ablation results under the inductive evaluation setup on the miniImageNet dataset using a ResNet-12 backbone.
Setup Acc (Ours) Cluster-FSL [ling2022semi]
5-way-1-shot
Transductive 84.51 77.81
Inductive 85.68 78.83
5-way-5-shot
Transductive 86.95 85.55
Inductive 87.23 85.64
8.5.3 Ablation and Sensitivity Analysis of CVOC and CST Hyperparameters
In this section, we ablate various hyperparameters used in the CVOC algorithm coupled with CST iterations in the finetuning stage. In the Figure  11, we present a heatmap analysis of the hyperparameters $w_{\text{intra}}$ and $w_{\text{inter}}$; while Table 5 reports results with and without incorporating intra and inter class terms. We defined the grid search over the following hyperparameter ranges:
  • $\epsilon\in\{1.0,2.0,3.0\}$
  • $\beta_{0}\in\{0.01,0.05,0.07\}$
  • $\gamma\in\{0.001,0.003,0.005,0.01\}$
  • $\alpha\in\{0.01,0.02,0.03\}$
The standard deviation of the downstream accuracy across the grid search was $\sigma=0.097$, indicating stable results. Refer to Figure 4 to see results with and without CST updates.
Figure 11: Heatmap showing the effect of $w_{\mathrm{intra}}$ and $w_{\mathrm{inter}}$ hyperparameters in the 5-way-1-shot setting and ResNet-12 backbone on miniImageNet dataset. Here, the values of $w_{\mathrm{intra}}$ vary row-wise, while the values of $w_{\mathrm{inter}}$ vary column-wise.
8.5.4 Ablation on the contributions of the Cluster Separation Tuner and Semantic Correction
@@
In this section, in order to isolate the contributions of the Cluster Separation Tuner (CST) and semantic correction, we disable the reconstruction-distance term and plug CST + semantic anchor directly into standard clustering methods. Here, we compare against the latest methods under the 5-way 1-shot and 5-shot tasks. Although we had to retune a few hyper-parameters - chiefly $s$ in the semantic anchor module and the number of CST iterations to fit this plug-in setup, the adjusted settings still yield competitive performance (Figure 8). We further incorporate Cluster-FSL equipped with Semantic Injection to clearly isolate its individual effect and to emphasize the modularity of the semantic injection mechanism.
8.5.5 Ablation on hyperparameter in Semantic anchor
@@
8.5.6 Ablation on Restricted Pseudo Labeling
At test time, we limit the augmentation of the support set by selecting only those unlabeled samples with the most confident pseudo-labels. Since pseudo-labels can be noisy, we sort unlabeled samples by their prediction entropy and pick the lowest-entropy examples, offering a simple, efficient strategy with minimal test-time overhead (see Figure 12). We observe that retaining 80% of the unlabeled samples yields the highest gains, highlighting the impact of confident pseudo-label selection. RPL performs well in every clean settings but may be less effective with distractor classes and very small unlabeled pools (<30) due to higher pseudo-label variance. However, with more than 50 unlabeled samples per class, it consistently improves accuracy even in distractor scenarios, while maintaining low inference overhead.
@@ 8.5.6 Ablation on Restricted Pseudo Labeling
8.5.7 Choice of embedding model
We evaluate two text encoders for generating class-description embeddings: a vanilla BERT-base model [devlin2019bert] and the CLIP text encoder [radford2021learning]. In both 5-way 1-shot and 5-way-5-shot settings, CLIP’s text features outperform BERT significantly (see Figure 13).
Potential reasons might be that CLIP’s text encoder is trained jointly with its image encoder on millions of image–caption pairs via a contrastive objective, grounding its representations in visual concepts. BERT, by contrast, learns only linguistic co-occurrence patterns and lacks alignment with vision. Consequently, CLIP’s text vectors naturally become a far more effective option for bridging semantic gaps in few-shot vision tasks than a standalone language model.
8.5.8 Robustness and Generalization of the Core Visual Framework
Even though our class description generation pipeline and semantic anchor network are major contributions, we conducted a series of rigorous experiments analyzing the performance of our method in the absence of any semantic guidance. These analyses demonstrate that the visual-only components alone establish a powerful and generalizable foundation for semi-supervised few-shot learning.
Ablation of Semantic Anchoring.
First, we evaluated our approach without the semantic anchoring module. As shown in the second row of our ablation study (Table 5), our method still outperforms recent state-of-the-art baselines.
Cross-Domain Generalization.
A critical test for any few-shot learning method is its ability to generalize to entirely unseen domains. We assessed the generalization capability of our core visual framework by training and fine-tuning it on miniImageNet [vinyals2016matching] and subsequently testing it on four diverse target datasets: CUB [hilliard2018few], EuroSAT  [helber2019eurosat], ISIC  [codella2019isic], and ChestX  [wang2017chestx], all without semantic anchoring. As detailed in  8.1.2, the framework demonstrates remarkably strong performance, validating that its learned structural priors are not confined to the source domain but transfer effectively to new visual concepts.
Qualitative Analysis of Class Activation.
Finally, we analyzed class activation maps (CAM) as described in  8.1.3. We generated heatmaps for images from the CUB dataset using a model trained only on miniImageNet. Compared to Cluster- FSL [ling2022semi], our framework’s heatmaps are more precisely localized on the object of interest, ignoring background clutter. This provides visual evidence that our methodology guide the model to focus on the true discriminative features of a class, leading to a more robust and interpretable representation space even in challenging cross-domain scenarios.
8.6 Additional Results on Modern Architectures (ViT & Swin)
To further demonstrate the generalizability of our method beyond standard CNN backbones, we evaluated SA-CVOC on modern Vision Transformer architectures, specifically ViT-Small [dosovitskiy2020image] and Swin-Tiny [liu2021swin]. As shown in Table 15, our method consistently outperforms the strong baseline Cluster-FSL [ling2022semi] across both architectures and settings on the miniImageNet dataset. The work of dosovitskiy2020image introduced the Vision Transformer, while liu2021swin proposed a hierarchical version. Our results show significant gains on Swin-Tiny, with an improvement of +6.7% in the 1-shot scenario and +1.09% in the 5-shot scenario, validating that our geometric regularization (CVOC) and semantic alignment modules are effective regardless of the underlying inductive bias (CNN vs. Transformer).
Table 15: Classification accuracy (%) on miniImageNet using Vision Transformer backbones.
Backbone Method 1-shot 5-shot
ViT-Small Cluster-FSL 77.90 86.15
Ours 84.20 87.21
Swin-Tiny Cluster-FSL 79.33 88.31
Ours 86.03 89.40
8.7 Limitations
Our approach assumes that both labeled and unlabeled data lie on a shared manifold, which is crucial for accurate pseudo-labeling. Our semantic anchoring technique relies on the choice of embedding model used to encode class descriptions. Consequently, this semantic anchor module may not work very well in complex datasets where the visual features of different classes are highly similar or where the visual data lacks distinctive semantic characteristics necessary for effective differentiation. Also, we compared our approach with a broader set of recent methods under transductive, semi-supervised settings. Although the network architecture and data splits may appear similar across works, we recognize that our results are only directly comparable to our own reproductions. This is because prior works differ significantly in key aspects such as network variants, training configurations, dataset versions, and preprocessing or even pretraining strategies.